OpenStudy (anonymous):

√12:(x5)*√6:(x) HELP PLEASE !!

OpenStudy (anonymous):

rewrite the equation with the equation tool | \/

OpenStudy (anonymous):

\[\sqrt{\frac{12}{x^5}}\times \sqrt{\frac{6}{x}}\]?

OpenStudy (anonymous):

\[6/x ^{3}\sqrt{2}\]

OpenStudy (anonymous):

sry, messed up on first answer

OpenStudy (anonymous):

π

OpenStudy (anonymous):

\[\sqrt{\frac{12*6}{x^5*x}}=\sqrt{\frac{72}{x^6}}\]

OpenStudy (anonymous):

\[\frac{6\sqrt{2}}{x^3}\]

OpenStudy (anonymous):

ty